OpenStudy (anonymous):

Prove this trig identity 2csc(x)tan(.5x)-1=tan^2(.5x)

OpenStudy (anonymous):

Here's a better way to show it \[2\csc x \tan \frac{ 1 }{ 2 }x=\tan ^{2}\frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

Oh sorry should have a -1 on the LHS

OpenStudy (jdoe0001):

hold the mayo

OpenStudy (jdoe0001):

ahh rats

OpenStudy (anonymous):

Thanks so much :) that's all I need

OpenStudy (jdoe0001):

\(\bf 2{\color{brown}{ csc(x) }} tan\left( \frac{x}{2} \right)-1=tan^2\left( \frac{x}{2}\right) \\ \quad \\ \quad \\ 2 {\color{brown}{\left[ csc\left( 2\cdot \frac{x}{2} \right) \right] }} tan\left( \frac{x}{2} \right)-1=tan^2\left( \frac{x}{2}\right) \\ \quad \\ 2 {\color{brown}{ \left[ \cfrac{1}{sin\left( 2\cdot \frac{x}{2} \right)} \right] }}tan\left( \frac{x}{2} \right)-1=tan^2\left( \frac{x}{2}\right) \\ \quad \\ \textit{now, using double identies for sine, we get} \\ \quad \\ 2 {\color{brown}{ \left[ \cfrac{1}{ 2sin\left(\frac{x}{2} \right)cos\left(\frac{x}{2}\right)} \right] }}tan\left( \frac{x}{2} \right)-1=tan^2\left( \frac{x}{2}\right) \\ \quad \\ \cancel{2}\left[ \cfrac{1}{ \cancel{2} \cancel{sin\left(\frac{x}{2} \right)} cos\left(\frac{x}{2}\right)} \right] {\color{brown}{ \cfrac{\cancel{sin\left(\frac{x}{2} \right)}}{cos\left(\frac{x}{2}\right)} }}-1=tan^2\left( \frac{x}{2}\right)\) there, see what you get check the pythagorean identities :)

OpenStudy (jdoe0001):

yw